A light work day, since we are still under a work stoppage for tower climbing. The crew spent about 4 hours attaching cables to the RMY sonics, assembling solar panel frames, assisting Cornell, and organizational tasks. The ops center WiFi access point is now in its final position working well. Gary got the ISS profiler site running again.
Tomorrow is a down day for our crew.